OpenStudy (anonymous):

Find all complex numbers z∈C with |z| < 2π such that -e^z =e^-z. Plot these points on the complex plane.

OpenStudy (anonymous):

so far I got that |z| < 2π is a circle (x^2 + y^2 < 4π^2) but Im not sure what to do about the -e^z = e^-z bit...

OpenStudy (anonymous):

@rational heya reckon you could help out?

OpenStudy (rational):

\(\dfrac{e^z + e^{-z}}{2} = \cosh (z)\)

OpenStudy (anonymous):

OHH yeahhh I forgot about that haha

OpenStudy (rational):

That may bot be useful hmm

OpenStudy (anonymous):

yea Im so lost

OpenStudy (rational):

Lets try factoring maybe.. \( -e^{z} =e^{-z} \) \( e^{z} +e^{-z}=0 \) \( e^{-z}(e^{2z} +1)=0 \)

OpenStudy (rational):

\(e^{-z}\) is never 0 so do we end up solving \(e^{2z}+1=0\) ?

OpenStudy (anonymous):

hmm how do we do that though?

OpenStudy (rational):

let \(e^z = t\)

OpenStudy (anonymous):

ohh righttt

OpenStudy (anonymous):

and then t = +-i?

OpenStudy (rational):

Yep!

OpenStudy (anonymous):

ohh so e^z = +-i

OpenStudy (rational):

Yes, look at the argand plane

OpenStudy (anonymous):

so z = +- π/2?

OpenStudy (rational):

Nope, lets look at the argand plane

OpenStudy (anonymous):

okay

OpenStudy (rational):



OpenStudy (anonymous):

is that the |z| < 2π part?

OpenStudy (rational):



OpenStudy (rational):

e^z = i e^(x+iy) = i e^x[cos(y) + isin(y)] = i compare real parts both sides : e^x cosy = 0 this gives x = 0, y = +- pi/2 therefore z = x + iy = 0+- ipi/2

OpenStudy (anonymous):

ohh okay that makes sense

OpenStudy (rational):

Notice, two solutions are +- ipi/2 not +- pi/2

OpenStudy (anonymous):

yup

OpenStudy (rational):

you should get two more solutions by solving e^z = -i

OpenStudy (anonymous):

would the two solutions just come out as +- πi/2 again?

OpenStudy (anonymous):

but how do you get +-3πi/2 ?

OpenStudy (rational):

Oh we're done actually, -pi/2 is same as 3pi/2

OpenStudy (anonymous):

ohhh I get it, you just add 2π because it still lies in the |z| < 2

OpenStudy (anonymous):

yeah lol thanks so much!! :D

OpenStudy (rational):

the four solutions are then +-pi/2 and +-3pi/2

OpenStudy (anonymous):

sorry I meant |z| < 2π *

OpenStudy (anonymous):

yup thank you!!!

OpenStudy (rational):

Np :)